OpenStudy (anonymous):

divide f(x) by d(x) and write a summary statement in the form indicated f(x)=x^4+5^3-2x^2+5x-3 d(x)=x^2+1

OpenStudy (whpalmer4):

shouldn't that be \[f(x) = x^4+5x^3-2x^2+5x-3\]?

OpenStudy (whpalmer4):

what is the first term of \(f(x)\) divided by the first term of \(d(x)\)? That will be the first term of your answer. Then multiply that term by \(d(x)\) and subtract the result from \(f(x)\). Now repeat the process with the new polynomial.

OpenStudy (anonymous):

yes, that's right, my bad. and i don't understand... @whpalmer4

OpenStudy (whpalmer4):

Okay. it's just like doing long division on numbers, except maybe easier! So, what is the first term of \(f(x)\)? What is the first term of \(d(x)\)?

OpenStudy (anonymous):

x^2 and x^2

OpenStudy (whpalmer4):

one of those is correct, but the other is not. \[f(x) = {\huge x^4} + 5x^3-2x^2+5x-3\]\[d(x) = {\huge x^2}+1\]

OpenStudy (anonymous):

yes! my bad!

OpenStudy (anonymous):

you just divide the first term? which means subtract the powers? x^2?

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Yes, \(x^4/x^2 = x^{4-2} = x^2\) So the first term of our answer is \(x^2\) Now we multiply that term by \(x^2+1\). \[x^2(x^2+1) = \]

OpenStudy (anonymous):

\[x^4 +x^2?\]

OpenStudy (whpalmer4):

Yes. Now we subtract that from \(f(x)\) to see what we have left. f(x) - (x^4+x^2) = x^4+5x^3-2x^2+5x-3 - (x^4+x^2) =

OpenStudy (whpalmer4):

\[f(x) - (x^4+x^2) = x^4+5x^3-2x^2+5x-3 - (x^4+x^2) =\]

OpenStudy (whpalmer4):

not my best day! need more coffee :-)

OpenStudy (anonymous):

\[5x^3-x^2+5x+3?\]

OpenStudy (anonymous):

and it's no problem!

OpenStudy (anonymous):

:)

OpenStudy (whpalmer4):

Hmm, close, but no cigar... What happened to the x^2 term?

OpenStudy (whpalmer4):

we're subtracting \((x^4+x^2)\) so that minus sign also is applied to the \(x^2\): \[x^4+5x^3-2x^2+5x-3-(x^4+x^2) = x^4+5x^3-2x^2+5x-3-x^4-x^2\]\[=x^4-x^4+5x^3-2x^2-x^2+5x-3 = 5x^3-3x^2+5x-3\] For this very reason, I prefer to multiply the result by -1 and add instead of trying to subtract. \[-1*x^2(x^2+1) = -x^4 - x^2\] \[f(x) + (-x^4 - x^2) = x^4+5x^3-2x^2+5x-3-x^4-x^2\]\[=5x^3-3x^2+5x-3\]

OpenStudy (whpalmer4):

Now we repeat the process with our new, smaller polynomial. What is the first term divided by the first term of \(d(x)\)?

OpenStudy (anonymous):

x^2

OpenStudy (whpalmer4):

\[5x^3/x^2 = x^2\]?

OpenStudy (anonymous):

5x!

OpenStudy (whpalmer4):

Good. The next term of our answer is \(5x\). What is \(5x*d(x)\)? That's what we'll subtract from \(5x^3-3x^2+5x-3\)

OpenStudy (anonymous):

5x^3+5x^2

OpenStudy (whpalmer4):

mmm....no. try again :-) \[5x*d(x) = 5x*(x^2+1) = \]

OpenStudy (anonymous):

5x^3+5x!

OpenStudy (whpalmer4):

Right. and when we remove that from what we have left of \(f(x)\) what do we get?

OpenStudy (anonymous):

would you be subtracting that from f(x)?

OpenStudy (whpalmer4):

no, from \(5x^3-3x^2+5x-3\)

OpenStudy (anonymous):

so -3x^2-3?

OpenStudy (whpalmer4):

Yes. Can you do the next step? Just like the first two...

OpenStudy (whpalmer4):

what is the first term of \(-3x^2-3\) divided by the first term of \(d(x)\)?